# Lambda-Skip Connections: the architectural component that prevents rank collapse

### v1

@@ 5.1 Comparison of Rank Collapse for Mamba Architecture with different values of $\lambda$
In Figure 2, we analyze how the rank collapse measure at the last layer varies with different skip connection strengths for both the Albert’s architecture (Lan et al. [2020]) and Mamba-2, keeping the full architecture (also LayerNorm). In both cases, we note the importance of tuning and choosing the correct value of $\lambda$ to avoid rank collapse. We see how a value of $\lambda$ very close to 0, as we expected, is detrimental and causes a very low value of $\mu$, even in the presence of LayerNorm, which is consistent with our findings in Theorem 4.3. Additionally, we see that a negative $\lambda$ leads to a higher $\mu$. Intuitively, the reason for this could be that a negative $\lambda$ can be interpreted as negative feedback in the system, hence helping stabilizing it. Note that even if in both cases our condition on $\lambda$ in Theorem 4.1 is too conservative (e.g. for Mamba-2 with $a=0.9$, $\lambda$ needs to satisfy$|\lambda|>937.924$), in practice much lower values of $\lambda$ are good enough to prevent rank collapse.
To investigate whether using $\lambda$ as a trainable parameter affects the learning performance, we train two transformer-based architectures and two SSM-based architectures on the image task of the long range arena (LRA) benchmark [Tay et al., 2020] and on a multi-query associative recall (MQAR) task [Arora et al., 2023]. The accuracy of all four architectures on both tasks are reported in Table 1 and the details on the experimental setup are provided in Appendix A.13. From the achieved accuracy we note that learning $\lambda$ does not affect the performance and even outperforms the models with fixed $\lambda$ in some cases.
@@ 6 Conclusion
This paper studies the phenomenon of rank collapse across general sequence model architectures, using the unifying framework from Dao and Gu [2024]. Our main contribution is the introduction of a lambda-skip connection, which adds a regulatory term to the standard skip connection component to control its strength. We provide sufficient conditions for this parameter under which rank collapse does not occur. We do this via a lower bound on the rank collapse measure, which holds for transformers, LTI SSMs, selective SSMs and more generally, all architectures that can be written in the framework from Dao and Gu [2024]. We further study the necessity of the lambda-skip connection by showing that when this component is ablated, selective SSMs also experience rank collapse even in the presence of LayerNorm. Moreover, we show that similarly to Transformers, selective SSMs can also be affected by doubly exponential rank collapse when both skip connections and LayerNorm are ablated. We suggest that the primary cause of this doubly exponential rank collapse is the input dependence of the $M^{(k)}$ matrix. We validate our theoretical analysis with simulation experiments and further analyze other components such as the Gating Mechanism, which we have found also play a role in rank collapse.
Limitations and Future Work. A limitation of this work is the exclusion of gating mechanisms from the theoretical analysis. Since gating mechanisms help to mitigate the effect of rank collapse, we expect their inclusion in the derivations would result in tighter upper bounds on the rank collapse measure for Mamba. Furthermore, one other line worth exploring is to consider the effect on lower bounds of MLPs on a theoretical standpoint instead of only considering skip connections and LayerNorm, which have not been considered in this paper. Finally, deriving a tighter lower bound that matches the results in our experiments is also a very promising avenue of research. Finally, Remark 4.1 offers intriguing insights from a control-theoretic perspective, as it suggests that the $\lambda$ parameter can be interpreted as a control gain. We believe that further theoretical analysis of the stability and performance associated with $\lambda$-skip connections could lead to a more principled design of learning architectures.
@@ A.4 Proof of Lower Bound of Rank Collapse Measure
In the following, we denote $V_{i}^{(k)}:=V_{:,i}^{(k)}$ and $\mu^{(k)}=\mu(Y^{(k)})$ for brevity.
$\tilde{\mu}^{{(k+1)}^{2}}$ $=\Big{|}\Big{|}\tilde{Y}^{(k+1)}-\frac{\bm{1}\bm{1}^{T}}{N}\tilde{Y}^{(k+1)}\Big{|}\Big{|}_{F}^{2}=\Big{|}\Big{|}(M^{(k)}+\lambda I)V^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}(M^{(k)}+\lambda I)V^{(k)}\Big{|}\Big{|}_{F}^{2}$
$=\lambda^{2}\Big{|}\Big{|}V^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}V^{(k)}\Big{|}\Big{|}_{F}^{2}+\sum_{i=1}^{d}\left\|M^{(k)}V_{i}^{(k)}-\bm{11^{T}}M^{(k)}\frac{V_{i}^{(k)}}{N}\right\|_{2}^{2}+$
$+\sum_{i=1}^{d}2\lambda\left(M^{(k)}V_{i}^{(k)}-\bm{11^{T}}M^{(k)}\frac{V_{i}^{(k)}}{N}\right)^{T}\left(V_{i}^{(k)}-\bm{11^{T}}\frac{V_{i}^{(k)}}{N}\right)$
$=\lambda^{2}\Big{|}\Big{|}V^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}V^{(k)}\Big{|}\Big{|}_{F}^{2}+\sum_{i=1}^{d}\left(M^{(k)}V_{i}^{(k)}-\bm{11^{T}}M^{(k)}\frac{V_{i}^{(k)}}{N}\right)^{T}\Bigg{(}\Bigg{.}M^{(k)}V_{i}^{(k)}-$
$-\bm{11^{T}}M^{(k)}\frac{V_{i}^{(k)}}{N}+2\lambda V_{i}^{(k)}-2\lambda\bm{11^{T}}\frac{V_{i}^{(k)}}{N}\Bigg{.}\Bigg{)}$
$=\lambda^{2}\Big{|}\Big{|}V^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}V^{(k)}\Big{|}\Big{|}_{F}^{2}+\sum_{i=1}^{d}\left(M^{(k)}V_{i}^{(k)}-\bm{11^{T}}M^{(k)}\frac{V_{i}^{(k)}}{N}\right)^{T}\Bigg{(}\Bigg{.}(M^{(k)}+2\lambda I)V_{i}^{(k)}-$
$\bm{11^{T}}(M^{(k)}+2\lambda I)\frac{V_{i}^{(k)}}{N}\Bigg{.}\Bigg{)}$
$=\lambda^{2}\Big{|}\Big{|}V^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}V^{(k)}\Big{|}\Big{|}_{F}^{2}+\sum_{i=1}^{d}\Bigg{(}\Bigg{.}V_{i}^{{(k)}^{T}}M^{{(k)}^{T}}(M^{(k)}+2\lambda I)V_{i}^{(k)}+$
$+\frac{V_{i}^{{(k)}^{T}}}{N}M^{{(k)}^{T}}\bm{11^{T}}(\bm{11^{T}})^{T}M^{(k)}\frac{V_{i}^{(k)}}{N}+\frac{V_{i}^{{(k)}^{T}}}{N}M^{{(k)}^{T}}\bm{11^{T}}(M^{(k)}+2\lambda I)V_{i}^{(k)}-$
$-V_{i}^{{(k)}^{T}}M^{{(k)}^{T}}\bm{11^{T}}(M^{(k)}+2\lambda I)\frac{V_{i}^{(k)}}{N}\Bigg{.}\Bigg{)}$
$=\lambda^{2}\Big{|}\Big{|}V^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}V^{(k)}\Big{|}\Big{|}_{F}^{2}+\sum_{i=1}^{d}V_{i}^{{(k)}^{T}}\Bigg{[}\Bigg{.}M^{{(k)}^{T}}(M^{(k)}+2\lambda I)+\frac{1}{N}M^{{(k)}^{T}}\bm{11^{T}}M^{(k)}-$
$-\frac{2}{N}M^{{(k)}^{T}}\bm{11^{T}}(M^{(k)}+2\lambda I)\Bigg{.}\Bigg{]}V_{i}^{(k)}$
$=\lambda^{2}\Big{|}\Big{|}V^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}V^{(k)}\Big{|}\Big{|}_{F}^{2}+\sum_{i=1}^{d}V_{i}^{{(k)}^{T}}\Bigg{[}M^{{(k)}^{T}}\left(I-\frac{1}{N}\bm{11^{T}}\right)M^{(k)}+$
$+2\lambda M^{{(k)}^{T}}\left(I-\frac{2}{N}\bm{11^{T}}\right)\Bigg{]}V_{i}^{(k)}$
$\geq\lambda^{2}\Big{|}\Big{|}V^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}V^{(k)}\Big{|}\Big{|}_{F}^{2}+\sum_{i=1}^{d}V_{i}^{{(k)}^{T}}\Bigg{[}2\lambda M^{{(k)}^{T}}\left(I-\frac{2}{N}\bm{11^{T}}\right)\Bigg{]}V_{i}^{(k)}$
where the last inequality follows from the fact that $M^{{(k)}^{T}}\left(I-\frac{1}{N}\bm{11^{T}}\right)M^{(k)}$ is Symmetric and Positive Semi-Definite.
@@
$\tilde{\mu}^{{(k+1)}^{2}}$ $\geq\lambda^{2}\Big{|}\Big{|}V^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}V^{(k)}\Big{|}\Big{|}_{F}^{2}-2\lambda\sum_{i=1}^{d}||V_{i}^{{(k)}}||^{2}\ ||M^{(k)}||\ \Bigg{|}\Bigg{|}\left(I-\frac{2}{N}\bm{11^{T}}\right)\Bigg{|}\Bigg{|}$
$\geq\lambda^{2}\Big{|}\Big{|}V^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}V^{(k)}\Big{|}\Big{|}_{F}^{2}-2\lambda\sum_{i=1}^{d}||V_{i}^{{(k)}}||^{2}\ ||M^{(k)}||$
$\geq\lambda^{2}\Big{|}\Big{|}V^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}V^{(k)}\Big{|}\Big{|}_{F}^{2}-2\lambda\sum_{i=1}^{d}||V_{i}^{{(k)}}||^{2}\ ||M^{(k)}||_{F}$
$=\lambda^{2}\Big{|}\Big{|}V^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}V^{(k)}\Big{|}\Big{|}_{F}^{2}-2\lambda||V^{{(k)}}||_{F}^{2}\ ||M^{(k)}||_{F}$
$\geq\lambda^{2}\Big{|}\Big{|}V^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}V^{(k)}\Big{|}\Big{|}_{F}^{2}-2\lambda N\sigma_{\text{max}}(C_{V}^{(k)})^{2}||M^{(k)}||_{F}$
where the last inequality follows since $||V^{(k)}||_{F}\leq||Y^{(k)}||_{F}\sigma_{\text{max}}(C_{V}^{(k)})$ and since we are using LayerNorm and hence every row of $Y^{{(k)}}$ has norm 1.
Furthermore, from Lemma A.1, we have that:
$\Big{|}\Big{|}V^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}V^{(k)}\Big{|}\Big{|}_{F}^{2}\geq\sigma_{\text{min}}(C_{V}^{(k)})^{2}\mu^{(k)^{2}}$
Hence, we have that:
$\tilde{\mu}^{{(k+1)}^{2}}$ $\geq c^{2}\lambda^{2}\mu^{(k)^{2}}-2\lambda NS^{2}||M^{(k)}||_{F}$
@@
Combining the previous two results, we then get:
$\mu^{{(k+1)}^{2}}\geq\frac{1}{\max_{i\in[N]}||\tilde{Y}^{(k+1)}_{i,:}||^{2}_{2}}\left(c^{2}\lambda^{2}\mu^{(k)^{2}}-2\lambda NS^{2}||M^{(k)}||_{F}\right)$
@@
$\max_{i\in[N]}||\tilde{Y}^{(k+1)}_{i,:}||^{2}_{2}$ $=\max_{i\in[N]}\sum_{j}\tilde{Y}_{ij}^{{(k+1)}^{2}}=\max_{i\in[N]}\sum_{j}\left(\sum_{l}\left(M^{(k)}+\lambda I\right)_{il}V^{{(k)}}_{lj}\right)^{2}$
$\leq\max_{i\in[N]}\sum_{l}\left(M^{(k)}+\lambda I\right)^{2}_{il}\sum_{j}\sum_{m}\underbrace{Y_{lm}^{(k)^{2}}}_{\leq 1}C_{V_{mj}}^{2}$
$\leq S^{2}(||M^{(k)}||_{F}+|\lambda|)^{2}$
Using the fact that $||M^{(k)}||_{F}\leq C_{M}$, we get:
$\mu^{{(k+1)}^{2}}\geq\frac{1}{S^{2}(C_{M}+|\lambda|)^{2}}\left(\lambda^{2}c^{2}\mu^{{(k)}^{2}}-2\lambda NS^{2}C_{M}\right)$
To guarantee that $\mu^{{(k+1)}^{2}}\geq a\mu^{{(k)}^{2}}$ for some $a<1$, we can simply lower bound the right hand side. Hence, we want that $\frac{1}{S^{2}(C_{M}+|\lambda|)^{2}}\left(\lambda^{2}c^{2}\mu^{{(k)}^{2}}-2\lambda NS^{2}C_{M}\right)\geq a\mu^{{(k)}^{2}}$.
@@
$\left(\lambda^{2}c^{2}-aS^{2}(C_{M}+|\lambda|)^{2}\right)\mu^{{(k)}^{2}}\geq 2\lambda NC_{M}$
To do so, we need to guarantee that $\lambda^{2}c^{2}-aS^{2}(C_{M}+|\lambda|)^{2}>0$. Hence, if $\lambda$ satisfies the condition above and we have that $\mu^{{(k)}^{2}}\geq\frac{2\lambda NS^{2}C_{M}}{\lambda^{2}c^{2}-aS^{2}(C_{M}+|\lambda|)^{2}}$, then $\mu^{{(k+1)}^{2}}\geq a\mu^{{(k)}^{2}}$. Note that in order to satisfy the condition on $\mu^{{(k)}^{2}}$ up until time $K$, we need to choose $\mu^{{(0)}^{2}}\geq\frac{1}{a^{K}}\frac{2\lambda NS^{2}C_{M}}{\lambda^{2}c^{2}-aS^{2}(C_{M}+|\lambda|)^{2}}$, which concludes the proof.

### v3

@@ 5.1 Comparison of Rank Collapse for Mamba Architecture with different values of $\lambda$
In Figure 2, we analyze how the rank collapse measure at the last layer varies with different skip connection strengths for both the Albert’s architecture (Lan et al. [2020]) and Mamba-2, keeping the full architecture (also LayerNorm). In both cases, we note the importance of tuning and choosing the correct value of $\lambda$ to avoid rank collapse. We see how a value of $\lambda$ very close to 0, as we expected, is detrimental and causes a very low value of $\mu$, even in the presence of LayerNorm, which is consistent with our findings in Theorem 4.3. Additionally, we see that a negative $\lambda$ leads to a higher $\mu$. Intuitively, the reason for this could be that a negative $\lambda$ can be interpreted as negative feedback in the system, hence helping stabilizing it. Note that even if in both cases our condition on $\lambda$ in Theorem 4.1 is too conservative, in practice much lower values of $\lambda$ are good enough to prevent rank collapse. In Figure 2 we note that as $|\lambda|$ increases, there is little variation in $\mu$. This can be explained using the lower bound proposed in Theorem 4.1. Specifically, as $|\lambda|\to\infty$, the parameter $a\to 1$. This implies that the rank collapse metric decreases much more slowly with the number of layers. In the limit, it does not decrease at all, remaining at or above the rank collapse measure at the input.
We investigate whether using $\lambda$ as a trainable parameter affects the learning performance. To this end, we train two transformer-based architectures and two SSM-based architectures on the image task of the long range arena (LRA) benchmark [Tay et al., 2020] and on a multi-query associative recall (MQAR) task [Arora et al., 2023]. The accuracy of all four architectures on both tasks are reported in Table 1, and the details on the experimental setup are provided in Appendix A.14. From our observations, we conclude that learning $\lambda$ does not affect the performance and even outperforms the models with fixed $\lambda$ in some cases.
@@ 6 Conclusion
This paper studies the phenomenon of rank collapse across general sequence model architectures, using the unifying framework from Ali et al. [2024] and Dao and Gu [2024]. Our main contribution is the introduction of a lambda-skip connection, which adds a regulatory term to the standard skip connection component to control its strength. We provide sufficient conditions for this parameter under which rank collapse does not occur. We do this via a lower bound on the rank collapse measure, which holds for transformers, LTI SSMs, selective SSMs and more generally, all architectures that can be written in the framework from Ali et al. [2024] and Dao and Gu [2024]. We further study the necessity of the lambda-skip connection by showing that when this component is ablated, selective SSMs also experience rank collapse even in the presence of LayerNorm. Moreover, we show that similarly to Transformers, selective SSMs can also be affected by doubly exponential rank collapse when both skip connections and LayerNorm are ablated. We suggest that the primary cause of this doubly exponential rank collapse is the input dependence of the $M^{(k)}$ matrix. We validate our theoretical analysis with simulation experiments and further analyze other components such as the Gating Mechanism, which we have found also play a role in rank collapse.
Limitations and Future Work. A limitation of this work is the exclusion of gating mechanisms from the theoretical analysis. Since gating mechanisms help to mitigate the effect of rank collapse, we expect their inclusion in the derivations would result in tighter upper bounds on the rank collapse measure for Mamba. Furthermore, one other line worth exploring is to consider the effect on lower bounds of MLPs in addition to skip connections and LayerNorm, Deriving a tighter lower bound that matches the results in our experiments is also a very promising avenue of research. Additionally, Remark 4.1 offers intriguing insights from a control-theoretic perspective, as it suggests that the $\lambda$ parameter can be interpreted as a control gain. We believe that further theoretical analysis of the stability and performance associated with $\lambda$-skip connections could lead to a more principled design of learning architectures. Finally, exploring alternative measures of rank collapse, like effective rank, and their correlation with our metric is a promising direction to expand on how our results generalize across metrics.
@@ A.4 Proof of Lower Bound of Rank Collapse Measure
In the following, we denote $V_{i}^{(k)}:=V_{:,i}^{(k)}$, $Y_{i}^{(k)}:=Y_{:,i}^{(k)}$ and $\mu^{(k)}=\mu(Y^{(k)})$ for brevity.
$\tilde{\mu}^{{(k+1)}^{2}}$ $=\Big{|}\Big{|}\tilde{Y}^{(k+1)}-\frac{\bm{1}\bm{1}^{T}}{N}\tilde{Y}^{(k+1)}\Big{|}\Big{|}_{F}^{2}=\Big{|}\Big{|}M^{(k)}V^{(k)}+\lambda Y^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}(M^{(k)}V^{(k)}+\lambda Y^{(k)})\Big{|}\Big{|}_{F}^{2}$
$=\lambda^{2}\Big{|}\Big{|}Y^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}Y^{(k)}\Big{|}\Big{|}_{F}^{2}+\sum_{i=1}^{d}\underbrace{\left\|M^{(k)}V_{i}^{(k)}-\bm{11^{T}}M^{(k)}\frac{V_{i}^{(k)}}{N}\right\|_{2}^{2}}_{\geq 0}+$
$+\sum_{i=1}^{d}2\lambda\left(M^{(k)}V_{i}^{(k)}-\bm{11^{T}}M^{(k)}\frac{V_{i}^{(k)}}{N}\right)^{T}\left(Y_{i}^{(k)}-\bm{11^{T}}\frac{Y_{i}^{(k)}}{N}\right)$
$\geq\lambda^{2}\Big{|}\Big{|}Y^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}Y^{(k)}\Big{|}\Big{|}_{F}^{2}+2\lambda\sum_{i=1}^{d}\left[\left(I-\frac{\bm{1}\bm{1}^{T}}{N}\right)M^{(k)}V_{i}^{(k)}\right]^{T}\left(I-\frac{\bm{1}\bm{1}^{T}}{N}\right)Y_{i}^{(k)}$
$=\lambda^{2}\Big{|}\Big{|}Y^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}Y^{(k)}\Big{|}\Big{|}_{F}^{2}+2\lambda\sum_{i=1}^{d}V_{i}^{(k)^{T}}M^{(k)^{T}}\left(I-\frac{\bm{1}\bm{1}^{T}}{N}\right)Y_{i}^{(k)}.$
@@
$\tilde{\mu}^{{(k+1)}^{2}}$ $\geq\lambda^{2}\Big{|}\Big{|}Y^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}Y^{(k)}\Big{|}\Big{|}_{F}^{2}-2\lambda\sum_{i=1}^{d}\ ||M^{(k)}||\ \Bigg{|}\Bigg{|}\left(I-\frac{1}{N}\bm{11^{T}}\right)\Bigg{|}\Bigg{|}V_{i}^{(k)^{T}}Y_{i}$
$=\lambda^{2}\Big{|}\Big{|}Y^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}Y^{(k)}\Big{|}\Big{|}_{F}^{2}-2\lambda||M^{(k)}||\sum_{i=1}^{d}\sum_{j=1}^{d}C_{V_{:,i}}^{T}Y^{(k)^{T}}Y_{i}^{(k)}$
$=\lambda^{2}\Big{|}\Big{|}Y^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}Y^{(k)}\Big{|}\Big{|}_{F}^{2}-2\lambda||M^{(k)}||\sum_{i=1}^{d}\sum_{j=1}^{d}C_{V_{ji}}Y_{j}^{(k)^{T}}Y_{i}^{(k)}$
$\geq\lambda^{2}\Big{|}\Big{|}Y^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}Y^{(k)}\Big{|}\Big{|}_{F}^{2}-2\lambda||M^{(k)}||\max_{i,j}|C_{V_{ji}}|\sum_{i=1}^{d}\sum_{j=1}^{d}Y_{j}^{(k)^{T}}Y_{i}^{(k)}$
$=\lambda^{2}\Big{|}\Big{|}Y^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}Y^{(k)}\Big{|}\Big{|}_{F}^{2}-2\lambda||M^{(k)}||\max_{i,j}|C_{V_{j,i}}|\sum_{i=1}^{d}\sum_{j=1}^{d}\sum_{l=1}^{N}Y_{lj}^{(k)}Y_{li}^{(k)}$
$=\lambda^{2}\Big{|}\Big{|}Y^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}Y^{(k)}\Big{|}\Big{|}_{F}^{2}-2\lambda||M^{(k)}||\max_{i,j}|C_{V_{j,i}}|\sum_{l=1}^{N}\left(\sum_{j=1}^{d}Y_{lj}^{(k)}\right)^{2}$
[…]
$\geq\lambda^{2}\Big{|}\Big{|}Y^{(k)}-\frac{\bm{1}\bm{1}^{T}}{N}Y^{(k)}\Big{|}\Big{|}_{F}^{2}-2\lambda||M^{(k)}||\max_{i,j}|C_{V_{j,i}}|Nd$
$\geq\lambda^{2}\mu^{(k)^{2}}-2\lambda||M^{(k)}||_{F}\ ||C_{V}||_{F}Nd$
where in (i) we used Cauchy-Schwartz inequality and the fact that we are using LayerNorm and hence every row of $Y^{{(k)}}$ has norm 1 and in the last step we used the definition of the rank collapse measure.
@@
Combining the previous two results and recalling that $C_{M}=\sup_{k}||M^{(k)}||_{F}$ and $S=\sup_{k}||C_{V}||_{F}$, we then get:
$\mu^{{(k+1)}^{2}}\geq\frac{1}{\max_{i\in[N]}||\tilde{Y}^{(k+1)}_{i,:}||^{2}_{2}}\left(\lambda^{2}\mu^{(k)^{2}}-2\lambda NdSC_{M}\right)$
@@
$\max_{i\in[N]}||\tilde{Y}^{(k+1)}_{i,:}||^{2}_{2}$ $=\max_{i\in[N]}\sum_{j}\tilde{Y}_{ij}^{{(k+1)}^{2}}=\max_{i\in[N]}\sum_{j}\left(\sum_{l}\left(M_{il}^{(k)}V^{{(k)}}_{lj}+\lambda Y_{lj}^{(k)}\right)\right)^{2}$
$\leq\max_{i\in[N]}\sum_{j}\sum_{l}\left(M^{(k)^{2}}_{il}V^{{(k)}^{2}}+\lambda^{2}_{lj}Y_{lj}^{(k)^{2}}+2\lambda M_{il}^{(k)}V^{{(k)}}_{lj}Y_{lj}^{(k)}\right)$
$=\max_{i\in[N]}\left(\sum_{l}M_{il}^{(k)^{2}}\sum_{j}\underbrace{V_{lj}^{(k)^{2}}}_{\leq S^{2}}+\lambda^{2}+2\lambda\sum_{l}M^{(k)}_{il}\sum_{j}V_{lj}^{(k)}Y_{lj}^{(k)}\right)$
$\leq\max_{i\in[N]}\left(S^{2}\underbrace{\sum_{l}M_{il}^{(k)^{2}}}_{\leq C_{M}^{2}}+\lambda^{2}+2\lambda\sum_{l}M^{(k)}_{il}\sum_{m}C_{V_{lm}}^{(k)}\underbrace{\sum_{j}Y_{mj}^{(k)}Y_{lj}^{(k)}}_{\leq 1}\right)$
$\leq\max_{i\in[N]}\left(S^{2}C_{M}^{2}+\lambda^{2}+2\lambda C_{M}S\right)$
$\leq(C_{M}S+\lambda)^{2}$
By putting everything together, we get:
$\mu^{{(k+1)}^{2}}\geq\frac{1}{(SC_{M}+|\lambda|)^{2}}\left(\lambda^{2}\mu^{{(k)}^{2}}-2\lambda NdSC_{M}\right)$
To guarantee that $\mu^{{(k+1)}^{2}}\geq a\mu^{{(k)}^{2}}$ for some $a<1$, we can simply lower bound the right hand side. Hence, we want that $\frac{1}{(SC_{M}+|\lambda|)^{2}}\left(\lambda^{2}\mu^{{(k)}^{2}}-2\lambda NdSC_{M}\right)\geq a\mu^{{(k)}^{2}}$.
@@
$\left(\lambda^{2}-a(SC_{M}+|\lambda|)^{2}\right)\mu^{{(k)}^{2}}\geq 2\lambda NdSC_{M}$
To do so, we need to guarantee that $\lambda^{2}-a(SC_{M}+|\lambda|)^{2}>0$. Hence, if $\lambda$ satisfies the condition above and we have that $\mu^{{(k)}^{2}}\geq\frac{2\lambda NdSC_{M}}{\lambda^{2}-a(SC_{M}+|\lambda|)^{2}}$ , then $\mu^{{(k+1)}^{2}}\geq a\mu^{{(k)}^{2}}$. Note that in order to satisfy the condition on $\mu^{{(k)}^{2}}$ up until time $K$, we need to choose $\mu^{{(0)}^{2}}\geq\frac{1}{a^{K}}\frac{2\lambda NdSC_{M}}{\lambda^{2}-a(SC_{M}+|\lambda|)^{2}}$, which concludes the proof.
In the following, we provide a brief discussion on the key variables of interest in the bound above, their typical values in practice, their relationship and implications on the bound, in particular focusing describing how it is possible to choose suitable values for $\lambda$. In the presented bound, the key variables of interest are $\lambda$, $a$ and $\mu(Y^{(k)})$. Specifically, we aim for $\mu(Y^{(k)})$ to be as far from 0 as possible, as values close to 0 indicate rank collapse. Furthermore, as we mentioned in the main text, the ideal value of $a$ would be 1 (since this would guarantee that the rank collapse metric is non decreasing over layers), although in order to satisfy Equation 7 this value cannot be chosen. In practice, the typical value of $\lambda$ is 1. The key relationship between $a$ and $\lambda$ is the following: in order to guarantee values of $a$ closer and closer to 1 (and hence to ensure higher values of the rank collapse metric at the final layer) we must choose larger values for $|\lambda|$. Additionally, $N$ represents the input sequence length, which varies based on the task. For example, $N$ might be on the order of tens for simple question answering tasks, but it could scale to hundreds or thousands when summarizing a long document. $d$ instead represents the embedding dimension, typically in the order of tens or hundreds. Regarding the selection of $\lambda$, we propose two possible approaches: treating it as a hyperparameter or making it learnable. In the first approach, $\lambda$ would be chosen through a standard hyperparameter optimization procedure, testing different values and evaluating their impact on both performance and the rank collapse measure. While effective, this method requires multiple training runs to identify the best value. In contrast, the second approach—making $\lambda$ learnable—offers significant advantages. It automates the process of finding an optimal $\lambda$, eliminating the need for manual hyperparameter tuning and requiring only a single training run. This is both more efficient and practical. For these reasons, we adopted the learnable $\lambda$ approach in our experiments, as illustrated in Table 1.